# Uncovering Treatment Effect Heterogeneity in Pragmatic Gerontology Trials

### v1

@@ 3.2 Principal stratification and estimands
We use the principal stratification framework frangakis2002principal; wang2024mixed to classify each individual by the pair of potential survival indicators $\big(S_{ij}(1),S_{ij}(0)\big)$. The four strata are:
@@
We adopt a monotonicity assumption tong2023bayesian; angrist1996identification stating that treatment does not reduce survival; equivalently, the harmed stratum is absent and $G_{ij}\in\{11,10,00\}$. Under this condition, observed data identify strata membership for two groups and leave two as mixtures: survivors observed under control must be always-survivors ($G_{ij}=11$), and non-survivors observed under treatment must be never-survivors ($G_{ij}=00$); by contrast, non-survivors under control comprise a mixture of protected and never-survivors ($G_{ij}\in\{10,00\}$), while survivors under treatment comprise a mixture of protected and always-survivors ($G_{ij}\in\{10,11\}$). In the WSD context, where telecare aims to improve well-being, monotonicity is substantively plausible.
@@
Let $X_{ij}$ denote baseline covariates (individual and cluster level). We also consider the Conditional SACE (CSACE),
@@ 3.3 Missing outcome and survival status
In the WSD trial, incompleteness arises from two sources: missing survival status at the end of follow-up and missing non-mortality outcomes (e.g., QoL measures). Let $R_{ij}^{S}$ and $R_{ij}^{Y}$ indicate whether survival status and the final outcome are observed, respectively ($1$ for observed and 0 for missing). When $S_{ij}=0$, the non-mortality outcome is undefined due to truncation by death, denoted $Y_{ij}=\star$. For reference, we summarize the feasible observation patterns and the recorded variables $\mathcal{O}_{ij}$ in Table 2; we will refer to each pattern by its label in bold.
@@
Death truncation $\{Z_{i},\ X_{ij},\ S_{ij}=0,\ R_{ij}^{S}=1\}$ (outcome undefined, $Y_{ij}=\star$)
@@
We formalize missingness using a nested MAR specification. Let $X_{ij}$ collect individual and cluster level covariates:
@@
In our analysis, for participants with observed survival but a missing final outcome we imputed the missing outcome using an arm-specific outcome model $Y_{ij}\mid(Z_{i},X_{ij},S_{ij}=1)$. When both survival status and the final outcome were unobserved we first sampling the principal stratum label $G_{ij}\in\{11,10,00\}$ from the nested Probit membership model introduced below; by monotonicity, $(G_{ij},Z_{i})$ maps to survival deterministically—survivors if $G_{ij}=11$ or if $G_{ij}=10$ and $Z_{i}=1$, and non-survivors otherwise. Those imputed as survivors then receive outcome imputation from the corresponding arm and stratum consistent model, whereas outcomes for non-survivors remain undefined ($Y_{ij}=\star$).
@@ 3.4 Likelihood and Bayesian principal stratification
We adopt a Bayesian principal stratification framework hirano2000assessing to infer principal stratum membership and to estimate SACE/CSACE. We can classify participants by observed treatment assignments and survival status at the final assessment:
@@
Under monotonicity, $O(1,0)$ are identified as never-survivors ($G_{ij}=00$) and $O(0,1)$ as always-survivors ($G_{ij}=11$); by contrast, $O(1,1)$ mixes $\{11,10\}$ and $O(0,0)$ mixes $\{10,00\}$. We need to do additional modeling to due with this mixture. We use $S^{obs}$ to denote the observed survival status for all individuals and $Y^{obs}$ to denote the collection of observed non-mortality outcomes for all individuals.
Let $p_{ij,g}=\Pr(G_{ij}=g\mid X_{ij},\theta_{S})$ denote the stratum membership probabilities (summing to one across $g$). For the outcome model, define the arm and stratum specific density evaluated at the observed outcome,
@@ 3.5 Model specification
We use a nested Probit construction to impute the unobserved principal strata labels $G_{ij}\in\{00,10,11\}$. chen2024bayesian Two latent Gaussian variables are augmented for each individual: $Q_{ij}$ first separates never-survivors $(00)$ from the survivor capable strata, and $W_{ij}$ then splits the survivor capable group into protected $(10)$ versus always-survivors $(11)$.
@@
For outcomes, we specify arm and stratum-specific Gaussian models (defined only when the potential outcome is meaningful in that stratum/arm):
$\{\,Y_{ij}(z)\mid G_{ij}=g,\ m_{g,z}(\bullet),\ \mathbf{X}_{Y,ij}\,\}\ \sim\ \mathcal{N}\!\big(m_{g,z}(\mathbf{X}_{g,z,ij}),\ \sigma^{2}_{g,z}\big),$
@@ 3.6 Integrating mixed-effects BART into principal stratification
To flexibly represent nonlinearities and higher order interactions in the conditional mean functions, we use Bayesian Additive Regression Trees (BART) chipman2010bart, which approximates a regression surface by a sum of shallow trees,
@@
$m_{g,z}(\mathbf{X}_{g,z,ij})=\sum_{k=1}^{K}g_{g,z,k}\!\big(\mathbf{X}_{g,z,ij};\mathcal{J}_{g,z,k},\mathcal{M}_{g,z,k}\big)+b_{i}^{(g,z)},\qquad(g,z)\in\{(11,0),(11,1),(10,1)\}.$
$b_{i}^{(g,z)}\sim\mathcal{N}(0,\sigma_{b,gz}^{2}).$
@@
Mixed-effects BART has several implementations, including stan4bart dorie2022stan, mixedbart spanbauer2021nonparametric, and hebart wundervald2022hierarchical. In simulations aligned with our design, mixedbart provided a favorable balance of accuracy and computational efficiency, and we therefore use it to estimate $m_{Q}$, $m_{W}$, and $m_{g,z}$. The mixedbart model is a general framework for incorporating hierarchical random effects into BART, originally developed for longitudinal clinical trials. Their approach supports both random intercepts and slopes, though in CRTs we adopt the special case of a cluster-level intercept.
@@ 3.7 Model estimation
A fully Bayesian joint model integrating imputation within the principal stratification sampler is also feasible but beyond the scope of this illustration. Missing baseline covariates and outcomes not attributable to death/serious deterioration were completed using multiple imputation rubin1978multiple to obtain a single analysis dataset. The covariate set used in the strata membership and outcome models comprised sex, age, ethnicity, highest education level, an indicator for living alone status, number of comorbid conditions, index of multiple deprivation (IMD) score , physical and mental health scores, and the baseline EQ-5D–VAS index.hirani2014effect
@@
We initialize the Markov chain by assigning principal–stratum labels using information identified by monotonicity: treated non-survivors are set to the never-survivor stratum and control survivors to the always-survivor stratum. For individuals in the remaining observed mixtures (treated survivors; control non-survivors), initial labels are drawn at random from the admissible strata given $(Z_{i},S_{ij})$. For units with unobserved survival, we impute their initial survival status by BART. Given these initial labels, initial estimates of outcome model $m_{g,z}(\cdot)$ can be obtained by fitting BART and initial estimate for $\sigma^{2}_{g,z}$ can be simultaneously obtained from this initial fit. For the strata–membership means $m_{Q}(\cdot)$ and $m_{W}(\cdot)$, we obtain starting values from parametric logistic regressions of the initial labels on covariates; latent normals $Q_{ij}$ and $W_{ij}$ are then drawn from the appropriate truncated Gaussian distributions implied by the initial labels. Hyperparameters for $\sigma_{g,z}^{2}$ use weakly informative inverse-gamma priors with shape and rate $a_{0}=b_{0}=0.001$ chen2024bayesian. For random intercept $\sigma_{b}^{2}$, we use the default prior in mxbart: hierarchical Inverse-Wishart with degree of freedom 3.
@@
Each iteration updates: (i) the outcome models $m_{11,1}$, $m_{11,0}$, and $m_{10,1}$ via mixed-effects BART on the current arm–stratum subsets; (ii) the residual variances $\{\sigma^{2}_{g,z}\}$ by conjugate Gaussian updates; (iii) the strata–membership means $m_{Q}$ and $m_{W}$ via mixed-effects BART regressions of the current latent normals $Q_{ij}$ and $W_{ij}$ on covariates; (iv) the latent $Q_{ij}$ and $W_{ij}$ from their truncated normal full conditionals; (v) principal-stratum labels $G_{ij}$ using their posterior probabilities that combine the current membership model and outcome likelihood; and (vi) missing data under the nested-MAR structure: survival is sampled from the strata model, outcomes are drawn from the appropriate arm– and stratum–specific outcome model when survival is (imputed) $1$, and outcomes are undefined by design when survival is 0. We used the default settings of mixedbart (mxbart) for Gaussian responses with a random intercept per cluster.

### v2

@@ 3.2 Principal stratification and estimands
We use the principal stratification framework 12; 37 to classify each individual by the pair of potential survival indicators $\big(S_{ij}(1),S_{ij}(0)\big)$. The four strata are:
@@
We adopt a monotonicity assumption 35; 2 stating that treatment does not reduce survival; equivalently, the harmed stratum is absent and $G_{ij}\in\{11,10,00\}$. Under this condition, observed data identify strata membership for two groups and leave two as mixtures: survivors observed under control must be always-survivors ($G_{ij}=11$), and non-survivors observed under treatment must be never-survivors ($G_{ij}=00$); by contrast, non-survivors under control comprise a mixture of protected and never-survivors ($G_{ij}\in\{10,00\}$), while survivors under treatment comprise a mixture of protected and always-survivors ($G_{ij}\in\{10,11\}$). In the WSD context, where telecare aims to improve well-being, monotonicity is substantively plausible.
@@
Let $X_{ij}$ denote baseline covariates (individual- and cluster-level). We also consider the Conditional SACE (CSACE),
@@ 3.3 Missing outcome and survival status
In the WSD trial, incompleteness arises from two sources: missing survival status at the end of follow-up and missing non-mortality outcomes (e.g., QoL measures). Let $R_{ij}^{S}$ and $R_{ij}^{Y}$ indicate whether survival status and the final outcome are observed, respectively ($1$ for observed and 0 for missing). When $S_{ij}=0$, the non-mortality outcome is undefined due to truncation by death, denoted $Y_{ij}^{\text{obs}}=\star$. For reference, we summarize the feasible observation patterns and the recorded variables $\mathcal{O}_{ij}$ in Table 2; we will refer to each pattern by its label in bold.
@@
Death truncation $\{Z_{i},\ X_{ij},\ S_{ij}=0,\ R_{ij}^{S}=1\}$ (outcome undefined, $Y_{ij}^{\text{obs}}=\star$)
@@
We formalize missingness using a nested MAR specification. Let $X_{ij}$ collect individual- and cluster-level covariates:
@@
In the WSD trial, Assumptions (1)-(2) encode a nested MAR structure: survival status is first subject to missingness through $R_{ij}^{S}$, and, conditional on being classified as a survivor with observed status ($S_{ij}=1$, $R_{ij}^{S}=1$), the final outcome is then subject to additional missingness through $R_{ij}^{Y}$. As described in Section 2, incomplete follow-up in WSD arose from three broad mechanisms: (i) truncation by death or serious clinical deterioration, which we treat as defining the nonsurvivor category for principal stratification; (ii) participants presumed alive but with unobserved QoL outcomes for non–health-related reasons (e.g., relocation or withdrawal from the study); and (iii) cases in which both survival status and QoL were unobserved due to administrative or logistical issues such as prolonged hospitalization, refusal to share records, or equipment problems.
Under cluster-level randomization, it is plausible that, conditional on the randomized treatment assignment $Z_{i}$ and the rich baseline covariates $X_{ij}$, the probability that survival status is recorded is driven primarily by these observed prognostic and administrative factors rather than by the unobserved potential survival or QoL outcomes themselves, which is exactly the content of Assumption (1). Likewise, among individuals with observed survival ($S_{ij}=1$, $R_{ij}^{S}=1$), missingness in the EQ-5D-VAS outcome is mainly attributable to patterns of loss to follow-up, non-response, or study withdrawal that are well captured by $(Z_{i},X_{ij})$, motivating Assumption (2) that $R_{ij}^{Y}$ is independent of the unobserved potential QoL outcomes given these observed quantities. In this sense, the nested MAR assumption provides a study-specific and practically interpretable approximation to the missing-data mechanism in the presence of truncation by death. To assess the impact of potential deviations from nested MAR, we further conduct simulation-based sensitivity analyses in which the missingness indicators are allowed to depend on latent prognostic variables beyond $(Z_{i},X_{ij})$ (Section 4.5).
In our analysis, we treat missing survival status and missing EQ-5D-VAS values as additional unknown quantities and impute them under the nested MAR assumptions in (1)-(2). The imputation scheme closely follows the four patterns in Table 2 and is embedded within our Bayesian estimation as a data–augmentation step.
First, for individuals in the “Survivor, outcome missing” pattern ($S_{ij}=1$, $R_{ij}^{S}=1$, $R_{ij}^{Y}=0$), we assume that, conditional on $(Z_{i},X_{ij})$, the outcome is missing at random. At each iteration of the Gibbs sampler, we draw the missing $Y_{ij}^{\text{mis}}$ from the arm- and stratum-specific outcome model
$Y_{ij}^{\text{mis}}\mid(Z_{i},X_{ij},S_{ij}=1,G_{ij}=g,\theta_{Y})\;\sim\;f_{Y}\!\big(\cdot\mid Z_{i},X_{ij},G_{ij}=g,\theta_{Y}\big),$
using the current draws of the model parameters $\theta_{Y}$ and the principal stratum label $G_{ij}$.
Second, for individuals in the “Status and outcome missing” pattern ($R_{ij}^{S}=0$), both $S_{ij}$ and $Y_{ij}$ are unobserved. For these subjects, we first sample the principal stratum membership $G_{ij}\in\{11,10,00\}$ from the nested Probit membership model described in Section 3.5, conditional on $(Z_{i},X_{ij})$ and the current latent variables $(Q_{ij},W_{ij})$. Under the monotonicity assumption, $(G_{ij},Z_{i})$ then determines survival deterministically: $S_{ij}=1$ if $G_{ij}=11$, or if $G_{ij}=10$ and $Z_{i}=1$, and $S_{ij}=0$ otherwise. For those with imputed $S_{ij}=1$, we draw $Y_{ij}^{\text{mis}}$ from the corresponding arm- and stratum-specific outcome model as above; for those with imputed $S_{ij}=0$, the non-mortality outcome remains undefined by design ($Y_{ij}^{\text{obs}}=\star$) and is not imputed.
@@ 3.4 Likelihood and Bayesian principal stratification
We adopt a Bayesian principal stratification framework 19 to infer principal stratum membership and to estimate SACE/CSACE. We can classify participants by observed treatment assignments and survival status at the final assessment:
@@
Under monotonicity, $O(1,0)$ are identified as never-survivors ($G_{ij}=00$) and $O(0,1)$ as always-survivors ($G_{ij}=11$); by contrast, $O(1,1)$ mixes $\{11,10\}$ and $O(0,0)$ mixes $\{10,00\}$. We need additional modeling to deal with this mixture. We use $S^{\text{obs}}$ to denote the observed survival status for all individuals and $Y^{\text{obs}}$ to denote the collection of observed non-mortality outcomes for all individuals.
Let $p_{ij,g}=\Pr(G_{ij}=g\mid X_{ij},\theta_{S})$ denote the stratum membership probabilities (summing to one across $g$). For the outcome model, define the arm- and stratum-specific density evaluated at the observed outcome,
@@ 3.5 Model specification
We use a nested Probit construction to impute the unobserved principal strata labels $G_{ij}\in\{00,10,11\}$. 6 Two latent Gaussian variables are augmented for each individual: $Q_{ij}$ first separates never-survivors $(00)$ from the survivor capable strata, and $W_{ij}$ then splits the survivor capable group into protected $(10)$ versus always-survivors $(11)$.
@@
For outcomes, we specify arm- and stratum-specific Gaussian models (defined only when the potential outcome is meaningful in that stratum/arm):
$\{\,Y_{ij}(z)\mid G_{ij}=g,\ m_{g,z}(\bullet),\ \mathbf{X}_{g,z,ij}\,\}\ \sim\ \mathcal{N}\!\big(m_{g,z}(\mathbf{X}_{g,z,ij}),\ \sigma^{2}_{g,z}\big),$
@@ 3.6 Integrating mixed-effects BART into principal stratification
To flexibly represent nonlinearities and higher order interactions in the conditional mean functions, we use Bayesian Additive Regression Trees (BART) 7, which approximates a regression surface by a sum of shallow trees,
@@
$m_{g,z}(\mathbf{X}_{g,z,ij})=\sum_{k=1}^{K}g_{g,z,k}\!\big(\mathbf{X}_{g,z,ij};\mathcal{J}_{g,z,k},\mathcal{M}_{g,z,k}\big)+b_{i}^{(g)},\qquad(g,z)\in\{(11,0),(11,1),(10,1)\}.$
$b_{i}^{(g)}\sim\mathcal{N}(0,\sigma_{b,g}^{2}).$
@@
Mixed-effects BART has several implementations, including stan4bart 9, mxbart 32, and hebart 39. In simulations aligned with our design, mxbart provided a favorable balance of accuracy and computational efficiency, and we therefore use it to estimate $m_{Q}$, $m_{W}$, and $m_{g,z}$. The mxbart model is a general framework for incorporating hierarchical random effects into BART, originally developed for longitudinal clinical trials. This approach supports both random intercepts and slopes, though in CRTs we adopt the special case of a cluster-level intercept.
@@ 3.7 Model estimation
A fully Bayesian joint model integrating imputation within the principal stratification sampler is also feasible but beyond the scope of this illustration. Missing baseline covariates were completed using multiple imputation 29 to obtain a single analysis dataset. The covariate set used in the strata membership and outcome models comprised sex, age, ethnicity, highest education level, an indicator for living alone status, number of comorbid conditions, index of multiple deprivation (IMD) score, physical and mental health scores, and the baseline EQ-5D-VAS index.18
@@
We initialize the Markov chain by assigning principal–stratum labels using information identified by monotonicity: treated non-survivors are set to the never-survivor stratum and control survivors to the always-survivor stratum. For individuals in the remaining observed mixtures (treated survivors; control non-survivors), initial labels are drawn at random from the admissible strata given $(Z_{i},S_{ij})$. For units with unobserved survival, we impute their initial survival status by BART. Given these initial labels, initial estimates of outcome model $m_{g,z}(\cdot)$ can be obtained by fitting BART and initial estimate for $\sigma^{2}_{g,z}$ can be simultaneously obtained from this initial fit. For the strata–membership means $m_{Q}(\cdot)$ and $m_{W}(\cdot)$, we obtain starting values from parametric logistic regressions of the initial labels on covariates; latent normals $Q_{ij}$ and $W_{ij}$ are then drawn from the appropriate truncated Gaussian distributions implied by the initial labels. Hyperparameters for $\sigma_{g,z}^{2}$ use weakly informative inverse-gamma priors with shape and rate $a_{0}=b_{0}=0.001$ 6. For random intercept $\sigma_{b}^{2}$, we use the default prior in mxbart: hierarchical Inverse-Wishart with degree of freedom 3.
@@
Each iteration updates: (i) the outcome models $m_{11,1}$, $m_{11,0}$, and $m_{10,1}$ via mixed-effects BART on the current arm–stratum subsets; (ii) the residual variances $\{\sigma^{2}_{g,z}\}$ by conjugate Gaussian updates; (iii) the strata–membership means $m_{Q}$ and $m_{W}$ via mixed-effects BART regressions of the current latent normals $Q_{ij}$ and $W_{ij}$ on covariates; (iv) the latent $Q_{ij}$ and $W_{ij}$ from their truncated normal full conditionals; (v) principal stratum labels $G_{ij}$ using their posterior probabilities that combine the current membership model and outcome likelihood; and (vi) missing data under the nested-MAR structure: survival is sampled from the strata model, outcomes are drawn from the appropriate arm– and stratum–specific outcome model when survival is (imputed) $1$, and outcomes are undefined by design when survival is 0. We used the default settings of mxbart for Gaussian responses with a random intercept per cluster.